# Personalized Reasoning: Just-in-time Personalization and Why LLMs Fail at It

### v1

@@ C.1 AIME
Oracle — Eval Summary Preference Weight Score Justification Comfort with Percentages 0.054 5 The response uses business-appropriate percentage terminology throughout, such as ’60% are girls,’ ’40% like to dance,’ and ’New percentage of girls: 58%.’ It assumes practical pe… Practical Logic Application 0.054 5 The response provides a clear and thorough logical explanation for each mathematical step. Step-by-step Structure Clarity 0.054 5 The response provides a very well-organized and sequential mathematical instruction. It begins with a clear visualization and analogy to the user’s interest (flower arrangement), … Analogy and Metaphor Use 0.043 5 The response uses multiple analogies and metaphors directly tied to the abstract concepts, making them concrete and understandable. Conceptual Insight Prioritization 0.043 5 The response skillfully integrates practical ’how’ (step-by-step calculations, algebraic setup, and solving the equation) with conceptual ’why’ (using the flower arrangement analo… Familiarity with Basic Algebra 0.043 5 The response uses concrete, step-by-step algebraic reasoning, starting with a relatable analogy (flower arrangement) and visualizing the problem in practical terms. Language Simplicity 0.043 5 The response uses very simple and accessible language throughout, such as ’Let me help you solve this problem in a way that feels comfortable and clear, just like arranging a deli… Patience and Reassurance 0.043 5 The response opens with a warm greeting (’Chào Lan!’) and immediately offers help in a way that feels ’comfortable and clear,’ using a gentle metaphor (’just like arranging a deli… Practical Application 0.043 5 The response makes a strong connection to Lan’s practical experience by using the flower shop analogy throughout, directly relating mathematical concepts to her business context. Real-World Connection 0.043 5 The response makes strong real-world connections by framing the mathematical problem in the context of the user’s flower shop, using the analogy of arranging a bouquet to represen… Narrative Framing 0.033 5 The response includes a strong narrative framing by using the flower shop and flower arrangement analogy throughout (e.g., ’just like arranging a delicate flower arrangement’, ’Im… Problem Complexity 0.033 5 The response breaks down the problem into clear, easy-to-understand steps, such as visualizing the scenario with a flower shop analogy, explicitly listing the initial percentages,… Explanation Depth 0.054 3 The response provides a step-by-step breakdown of the problem, including setting up variables, applying percentages, and solving the equation. Hands-on Engagement Opportunity 0.054 3 The response provides some hands-on opportunities by inviting the user to visualize the problem and follow step-by-step calculations (e.g., ’Let’s Set Up Our Flower Shop Calculati… Preferred Format 0.054 3 The response includes visual components, such as the flower shop analogy and flower emojis (), and textual explanations, such as step-by-step breakdowns and algebraic calculat… Trial-and-Error Support 0.054 3 The response provides some opportunities for trial-and-error by breaking down the problem into steps and inviting the user to ask for further explanation (’Would you like me to ex… Visual Learning Style 0.054 3 The response attempts to use visual metaphors, such as comparing students to flowers in a bouquet and referencing a ’flower arrangement’ to make the abstract mathematical concepts… Gardening Growth Analogies 0.043 3 The response uses some gardening and floral design analogies, such as ’arranging a delicate flower arrangement,’ ’each student is like a unique flower in a bouquet,’ and reference… Interactivity Level 0.043 3 The response provides some opportunities for engagement, such as inviting the user to request further explanation (’Would you like me to explain any part in more detail, like how …
(cont’d) Oracle — Eval Summary Preference Weight Score Justification Patience with Repetition 0.043 3 The response provides a step-by-step breakdown of the problem and uses a flower arrangement analogy to aid understanding, which demonstrates some patience and an attempt to explai… Collaborative Brainstorming 0.033 3 The response offers a friendly, personalized explanation and uses analogies that could foster a sense of community (e.g., ’Imagine your flower shop, where each student is like a u… Traditional Learning Methods 0.022 3 The response incorporates traditional wisdom and cultural perspectives by using the metaphor of arranging a delicate flower arrangement and referencing a flower shop, which are cu… Cultural Context 0.011 3 The response uses respectful and gentle language, such as ’Chào Lan!’ and analogies like ’arranging a delicate flower arrangement,’ which are generally positive and could be inter…

### v2

@@ C.1 AIME
Oracle — Eval Summary Preference Weight Score Justification Comfort with Percentages 0.054 5 The response uses business-appropriate percentage terminology throughout, such as ’60% are girls,’ ’40% like to dance,’ and ’New percentage of girls: 58%.’ It assumes practical pe… Practical Logic Application 0.054 5 The response provides a clear and thorough logical explanation for each mathematical step. Step-by-step Structure Clarity 0.054 5 The response provides a very well-organized and sequential mathematical instruction. It begins with a clear visualization and analogy to the user’s interest (flower arrangement), … Analogy and Metaphor Use 0.043 5 The response uses multiple analogies and metaphors directly tied to the abstract concepts, making them concrete and understandable. Conceptual Insight Prioritization 0.043 5 The response skillfully integrates practical ’how’ (step-by-step calculations, algebraic setup, and solving the equation) with conceptual ’why’ (using the flower arrangement analo… Familiarity with Basic Algebra 0.043 5 The response uses concrete, step-by-step algebraic reasoning, starting with a relatable analogy (flower arrangement) and visualizing the problem in practical terms. Language Simplicity 0.043 5 The response uses very simple and accessible language throughout, such as ’Let me help you solve this problem in a way that feels comfortable and clear, just like arranging a deli…
(cont’d) Oracle — Eval Summary Preference Weight Score Justification Patience and Reassurance 0.043 5 The response opens with a warm greeting (’Chào Lan!’) and immediately offers help in a way that feels ’comfortable and clear,’ using a gentle metaphor (’just like arranging a deli… Practical Application 0.043 5 The response makes a strong connection to Lan’s practical experience by using the flower shop analogy throughout, directly relating mathematical concepts to her business context. Real-World Connection 0.043 5 The response makes strong real-world connections by framing the mathematical problem in the context of the user’s flower shop, using the analogy of arranging a bouquet to represen… Narrative Framing 0.033 5 The response includes a strong narrative framing by using the flower shop and flower arrangement analogy throughout (e.g., ’just like arranging a delicate flower arrangement’, ’Im… Problem Complexity 0.033 5 The response breaks down the problem into clear, easy-to-understand steps, such as visualizing the scenario with a flower shop analogy, explicitly listing the initial percentages,… Explanation Depth 0.054 3 The response provides a step-by-step breakdown of the problem, including setting up variables, applying percentages, and solving the equation. Hands-on Engagement Opportunity 0.054 3 The response provides some hands-on opportunities by inviting the user to visualize the problem and follow step-by-step calculations (e.g., ’Let’s Set Up Our Flower Shop Calculati… Preferred Format 0.054 3 The response includes visual components, such as the flower shop analogy and flower emojis (), and textual explanations, such as step-by-step breakdowns and algebraic calculat… Trial-and-Error Support 0.054 3 The response provides some opportunities for trial-and-error by breaking down the problem into steps and inviting the user to ask for further explanation (’Would you like me to ex… Visual Learning Style 0.054 3 The response attempts to use visual metaphors, such as comparing students to flowers in a bouquet and referencing a ’flower arrangement’ to make the abstract mathematical concepts… Gardening Growth Analogies 0.043 3 The response uses some gardening and floral design analogies, such as ’arranging a delicate flower arrangement,’ ’each student is like a unique flower in a bouquet,’ and reference… Interactivity Level 0.043 3 The response provides some opportunities for engagement, such as inviting the user to request further explanation (’Would you like me to explain any part in more detail, like how … Patience with Repetition 0.043 3 The response provides a step-by-step breakdown of the problem and uses a flower arrangement analogy to aid understanding, which demonstrates some patience and an attempt to explai… Collaborative Brainstorming 0.033 3 The response offers a friendly, personalized explanation and uses analogies that could foster a sense of community (e.g., ’Imagine your flower shop, where each student is like a u… Traditional Learning Methods 0.022 3 The response incorporates traditional wisdom and cultural perspectives by using the metaphor of arranging a delicate flower arrangement and referencing a flower shop, which are cu… Cultural Context 0.011 3 The response uses respectful and gentle language, such as ’Chào Lan!’ and analogies like ’arranging a delicate flower arrangement,’ which are generally positive and could be inter…
@@ C.3 SocialIQA
Appendix D User Simulator Prompts
Passive User
Passive User Prompt You are role-playing as a passive and economical human user who is interacting with an AI assistant. Your goal is to generate a realistic response that reflects a user who minimizes their effort. YOUR PERSONA {persona_profile} CHAT HISTORY {chat_history} YOUR TASK The assistant has just said something. As a passive user, your goal is to answer the question factually but with the minimum information necessary. CORE INSTRUCTIONS 1. Stay in Character: You are a passive user. You are not proactive or collaborative. You answer what is asked and then stop. 2. Provide Factual, Atomic Answers: When asked a question, consult your persona and provide a direct, factual answer. Do not add extra context, opinions, or follow-up questions. Your answers should be “atomic”—the smallest possible unit of factual information that directly answers the question. 3. NEVER Volunteer Information: Do not offer information that wasn’t explicitly requested. The assistant must ask follow-up questions to get more details. 4. Reveal Preferences Only if Directly Asked: Only state a preference if the assistant asks a direct question like, “Would you prefer a simple or detailed explanation?” Your answer should be minimal (e.g., “I prefer a simple one.”). OUTPUT FORMAT Provide your response as a clean JSON object with two keys: "thought": A brief thought process explaining your minimal reasoning. For example: “The assistant asked about my background. The most direct atomic fact is my profession. I will state that and nothing else.” or “This looks correct. I’m done.” "response": The user’s actual, direct, and minimal response to the assistant.
Collaborative User
Collaborative User Prompt You are role-playing as a collaborative and proactive human user who is interacting with an AI assistant. Your goal is to generate a realistic and helpful response that reflects this personality. YOUR PERSONA {persona_profile} CHAT HISTORY {chat_history} YOUR TASK The assistant has just said something. As a collaborative user, your goal is to actively help the assistant give you the best possible answer. CORE INSTRUCTIONS 1. Stay in Character: You are a helpful and engaged user. You are willing to provide details to get to the solution faster. 2. Give Detailed, Contextual Answers: When the assistant asks a question, provide a comprehensive answer and include any extra context that might be useful. 3. Volunteer Information Freely: If you think of a detail that could help the assistant, offer it without being asked. Feel free to share your preferences, current understanding, and goals. 4. Ask Clarifying Questions: If the assistant’s response is good but could be better, ask follow-up questions to refine the answer. 5. Check for Satisfaction: If the assistant’s response has fully solved your problem and met your preferences, your response should be the termination signal. 6. Termination Signal: When you are completely satisfied, respond with ONLY the termination signal: {terminal_signal} OUTPUT FORMAT Provide your response as a clean JSON object with two keys: "thought": A brief thought process explaining your helpful reasoning. For example: “The assistant asked about my background. I’$l$ not only answer but also mention my learning goal to give them more context.” or “This is a good start, but I’$l$ ask a follow-up question to get more detail on the second step.” "response": The user’s actual, detailed response to the assistant.
Appendix E Assistant Prompts for Different Modes
Oracle Mode
Persona Known System Prompt You are a helpful assistant trying to generate a personalized explaination to the problem the user asks you. The user has the following Persona: {persona_profile} Respond in a way that aligns with these preferences. You will be evaluated on how well your explaination aligns with these preferences, so that is your primary goal.
Discovery Mode (Dynamic)
Infer Persona System Prompt (Dynamic Version) You are a helpful assistant. Before attempting to solve the user’s task or answer their question, first aim to understand the user’s background, goals, and preferences. Start by asking one or two clarifying questions that will help tailor your response to their needs. Do not proceed with a full answer until you have enough information to personalize it effectively. Your tone should remain natural, curious, and respectful—avoid interrogating the user, but try to guide the conversation to learn more about them. At each turn, you need to decide whether to ask the user for more information or to proceed with the answering the question. If you decide to ask for more information, you must ask one or two clarifying questions that will help tailor your response to their needs. Do not proceed with a full answer until you have enough information to personalize it effectively. Your response must be in the following format for all turns: ###ACTION###: ’ask_question’ or ’final_answer’ ###RESPONSE###: [clarifying question] or [full answer]
Discovery Mode (Fixed Number of Questions)
Infer Persona System Prompt (Fixed Questions Version) You are a helpful assistant. You must ask exactly {fixed_num_questions} clarifying questions to understand the user’s background, goals, and preferences before providing a final answer. This is turn {current_turn} of {fixed_num_questions}. Ask one or two clarifying questions that will help tailor your response to their needs. Your tone should remain natural, curious, and respectful—avoid interrogating the user, but try to guide the conversation to learn more about them. Your response must be in the following format: ###ACTION###: ’ask_question’ ###RESPONSE###: [clarifying question]
Baseline Mode
No Prompt System Prompt (No system prompt provided)
Appendix F Benchmark Generation Prompts
Preference Dimension Sampler
Preference Dimension Sampling Prompt Generate {min_dimensions_per_problem} preference dimensions for how someone would want this {domain} explained. PROBLEM: {problem_text} REQUIRED DIMENSION TYPES: 1. Expertise/Comfort Dimensions (2-3 dimensions): Identify the key skills/concepts needed for this specific problem and create dimensions about comfort level with those skills. Examples: ‘Comfort with Recursive Thinking’ (for recursion problems) ‘Familiarity with Inequality Manipulation’ (for inequality problems) ‘Experience with Algorithm Complexity’ (for efficiency problems) Only include expertise dimensions for skills that are ACTUALLY needed for this problem. 2. Personal Dimensions (remaining dimensions): Think beyond generic categories. Focus on authentic, personal aspects that would make someone think ‘this person really gets how I need to learn.’ CREATIVE REFLECTION PROCESS: What unique ways might someone want information delivered based on their life experiences? How might someone’s profession, personality, or background create unusual learning preferences? What would make someone think ‘Yes, this explanation style is perfect for me’? Each dimension should be: Specific and actionable (something that could guide how to explain) Personally meaningful (something someone would actually care about) Grounded in authentic human needs and experiences NOW, generate these dimensions from the perspective of this specific person: PERSONA PROFILE: Name: {persona_name} Summary: {minimal_necessary_description} Age: {age}, Occupation: {occupation} Location: {location} Hobbies: {hobbies} Domain expertise: {domain_expertise} LEARNING PROFILE: Knowledge level: {knowledge_level} Learning style: {learning_style} Problem-solving: {problem_solving} Confidence: {confidence} Anxiety triggers: {anxiety_triggers} Motivation: {motivation} PERSONALITY: Openness: {openness_description} Conscientiousness: {conscientiousness_description} Social style: {extraversion_description} BACKSTORY: {backstory} INSTRUCTIONS: 1. First, analyze the problem to identify 2-3 key skills/concepts needed to solve it 2. Create expertise dimensions for those specific skills based on this person’s background 3. Then create personal dimensions based on their unique characteristics Based on this person’s unique background, personality, experiences, and current situation, what specific aspects of an explanation would matter most to THEM? Generate dimensions that capture the nuanced, personal ways THIS person would want the {domain} explained. Think about how their professional experience, hobbies, personality, learning challenges, and life story create unique preferences that wouldn’t apply to everyone. 4. Ensure the total is exactly {min_dimensions_per_problem} dimensions NAMING REQUIREMENTS: Use concise, direct names (e.g., ‘Visual Learning Style’, ‘Technical Depth’) DO NOT use prefixes like ‘Preference for’, ‘Desire for’, ‘Need for’ DO NOT use redundant words like ‘Preference’, ‘Style’, ‘Approach’ unless essential Make names specific and actionable (what the preference controls) Format your response as a clean JSON object: { "dimensions": [ { "name": "(example) Technical Depth", "description": "Why this matters and how it affects learning experience", "value_range": "1-5 scale or categorical options", "type": "expertise" or "personal" } ] }
Preference Instantiation
Preference Instantiation Prompt You are an expert in personalization systems and educational psychology. Given a persona and relevant preference dimensions, generate this persona’s specific preference values. PERSONA SUMMARY: Name: {persona_name} Background: {minimal_necessary_description} Key Learning Characteristics: Cognitive Style: {learning_style} Problem Solving: {problem_solving} Confidence: {confidence} Personality Traits: Openness={openness}, Conscientiousness={conscientiousness}, Extraversion={extraversion}, Agreeableness={agreeableness}, Neuroticism={neuroticism} Domain Expertise: {domain_expertise} CURRENT PROBLEM: {problem_text} RELEVANT PREFERENCE DIMENSIONS: {dimension_list} [If persona has existing preferences] PERSONA’S EXISTING PREFERENCES FROM PREVIOUS PROBLEMS: {existing_preferences} CRITICAL INSTRUCTIONS: 1. Consistency: If this persona has existing preferences for similar dimensions, maintain consistency with their established patterns 2. Transferability: Consider how preferences might transfer between problem types. For example: Visual preferences might transfer well between geometry and coding Communication tone preferences should be highly consistent across problems Detail level preferences might vary based on problem complexity 3. Persona-Specific: Ground all preferences in this persona’s characteristics and background 4. Problem-Relevant: Adjust importance and some values based on this specific problem type 5. Justification: Explain how you considered existing preferences and problem context For each dimension, provide: 1. Value: Specific value that fits this persona’s profile and is consistent with existing preferences 2. Local Importance: 1-5 scale of how much this persona cares about this aspect for THIS specific problem. Make sure to spread out the range of importance; for example you cannot say all dimensions are important. 3. Justification: Detailed explanation of your reasoning, including reference to existing preferences if relevant Format as JSON: { "preferences": { "Dimension Name": { "name": "Copy the name of the preference dimension from the previous step", "description": "Copy the description of the preference dimension from the previous step", "value_range": "Copy the value range of the preference dimension from the previous step", "value": "specific value or number", "local_importance": 1-5, "justification": "Detailed reasoning considering persona characteristics, existing preferences, and problem context", "type": "expertise" or "personal" } } }
Evaluation Rubric Generator (Expertise Criteria)
Expertise Rubric Generation Prompt You are creating evaluation criteria for EXPERTISE/COMFORT preferences—how well an AI response matches a user’s technical skill level and comfort with specific concepts. PROBLEM TO BE ANSWERED: Type: {problem_type} Problem: {problem_text} [If answer is available] Correct answer: {solution_context} PERSONA CONTEXT: User: {persona_name} - {minimal_necessary_description} Background: {backstory} EXPERTISE/COMFORT PREFERENCES: {expertise_preferences_with_weights} EXPERTISE RUBRIC GUIDELINES: Focus on TECHNICAL APPROPRIATENESS: Does the response match the user’s stated comfort/skill level? Evaluate CONCEPT COMPLEXITY: Are concepts presented at the right level of sophistication? Assess TERMINOLOGY USAGE: Is technical language appropriate for their experience level? Check ASSUMPTION LEVELS: Does the response assume the right background knowledge? CRITICAL: Create exactly ONE criterion for each expertise preference listed above. Each criterion must use the EXACT preference name. For each expertise preference, create criteria that measure how well the AI response calibrates its technical level to match the user’s capabilities for this specific {problem_type}. EXAMPLE EXPERTISE CRITERION: { "preference": "Comfort with Linear Algebra", "description": "How well the response matches the user’s intermediate comfort level (value: 3) with linear algebra when explaining matrix operations in this problem", "weight": 0.25, "levels": [ {"score": 1, "description": "Uses linear algebra concepts far above (graduate level) or below (basic arithmetic) the user’s intermediate level, making explanation inaccessible or patronizing"}, {"score": 3, "description": "Mostly appropriate for intermediate level but some inconsistencies---occasionally too advanced or too basic for stated comfort level"}, {"score": 5, "description": "Perfectly calibrated to intermediate linear algebra comfort---uses appropriate terminology, assumes right background knowledge, explains concepts at ideal complexity level"} ] } Format as JSON: { "criteria": [ { "preference": "Exact preference name", "description": "How well the response matches the user’s technical level for this specific concept in this problem", "weight": 0.XX, "levels": [ {"score": 1, "description": "Technical level far above or below user’s stated comfort level"}, {"score": 3, "description": "Mostly appropriate technical level but some inconsistencies"}, {"score": 5, "description": "Perfectly calibrated to user’s stated comfort level"} ] } ] }
Evaluation Rubric Generator (Personal Criteria)
Personal Rubric Generation Prompt You are creating evaluation criteria for PERSONAL preferences—how well an AI response adapts to a user’s unique learning style, communication preferences, and personal context. PROBLEM TO BE ANSWERED: Type: {problem_type} Problem: {problem_text} [If answer is available] Correct answer: {solution_context} PERSONA CONTEXT: User: {persona_name} - {minimal_necessary_description} Background: {backstory} PERSONAL PREFERENCES: {personal_preferences_with_weights} PERSONAL RUBRIC GUIDELINES: Focus on LEARNING STYLE ADAPTATION: Does the response match how this person prefers to learn? Evaluate COMMUNICATION STYLE: Is the tone, formality, and interaction style appropriate? Assess CONTEXTUAL RELEVANCE: Are examples, analogies, and references meaningful to this person? Check MOTIVATIONAL ELEMENTS: Does the response connect to what drives this person? Consider this person’s background: {occupation}, hobbies: {hobbies}, and learning style: {learning_style} CRITICAL: Create exactly ONE criterion for each personal preference listed above. Each criterion must use the EXACT preference name. For each personal preference, create criteria that measure how well the AI response adapts to this specific person’s unique learning needs and preferences for this {problem_type}. EXAMPLE PERSONAL CRITERION: { "preference": "Engineering Examples from Work Experience", "description": "How well the response incorporates relevant engineering examples that connect to this user’s mechanical engineering background when explaining this physics problem", "weight": 0.20, "levels": [ {"score": 1, "description": "No engineering examples provided, or examples from completely irrelevant fields that don’t connect to user’s mechanical engineering experience"}, {"score": 3, "description": "Includes some engineering examples but they’re generic or don’t clearly connect to the physics concepts in this specific problem"}, {"score": 5, "description": "Provides multiple relevant mechanical engineering examples that directly illuminate the physics concepts and clearly resonate with the user’s professional experience"} ] } Format as JSON: { "criteria": [ { "preference": "Exact preference name", "description": "How well the response adapts to this user’s specific personal learning preference for this problem", "weight": 0.XX, "levels": [ {"score": 1, "description": "No adaptation to personal preference or completely inappropriate approach"}, {"score": 3, "description": "Some adaptation to personal preference but generic or inconsistent"}, {"score": 5, "description": "Excellent adaptation that perfectly matches the user’s personal preference"} ] } ] }
F.1 LLM Judge Scoring Prompt
The judge evaluates each rubric criterion independently using the following prompts.
System Message You evaluate responses against specific criteria. Output valid JSON.
User Prompt You are an expert evaluation specialist using a standardized rubric to assess personalized responses. Provide a precise, evidence-based assessment of how well this response meets the specified criterion. CRITERION TO EVALUATE:
{criterion name}
Description: {criterion description}
PERFORMANCE LEVELS:
{JSON list of 5 rubric levels with descriptions}
RESPONSE TO EVALUATE:
{model’s final response}
USER CONTEXT:
Preference value: {user’s preferred value on 1--5 scale}
Context: {justification for why this level suits the user}
EVALUATION INSTRUCTIONS: 1. Carefully compare the response against each performance level description. 2. Identify specific evidence in the response that matches level descriptions. 3. Determine the exact score (1–5) that best represents the response quality for this criterion. 4. Provide a detailed justification referencing specific elements of the response. 5. Be objective and consistent—apply the same standards across all evaluations. 6. Consider the user’s stated preferences. 7. Provide precise reasoning that another evaluator could follow to reach the same conclusion. Format your evaluation as a JSON object: {"score": X (integer between 1 and 5), "justification": "..."}
Rubric-Level Scoring.
The rubric performance levels are personalized at benchmark-construction time to each user’s preferred value $v_{j}$. For instance, for a user who prefers “Terminology Complexity” =2=2 (simple language), the rubric levels are defined as:
Example: Personalized Rubric Levels for Terminology Complexity ($v_{j}=2$) Score 1: Uses highly technical jargon far above the user’s preferred simple language level, making the explanation inaccessible. Score 3: Mostly uses accessible language but occasionally introduces unnecessary technical terms, inconsistently matching the user’s preference for simplicity. Score 5: Uses plain, accessible terminology throughout and avoids unnecessary jargon, matching the user’s preference for simple language.
The judge receives these pre-personalized level descriptions and matches the response against them. The grading function $g_{j}(r,v_{j})$ (Eq. 1) therefore measures how well the response matches the user’s stated preference level for each attribute.